OpenStudy (anonymous):

Help!! Giving out medals ;) Find the common ratio of the sequence -164, -82, -41, -20.5,.. A. 2 B. -82 C. 1/2 D. 82

OpenStudy (dtan5457):

You asked this question before... what does the pattern look like to you?

OpenStudy (dtan5457):

-164, -82, -41, -20.5,.. does the numbers look like they are getting..i don't know..halved?

OpenStudy (dtan5457):

@gxldmouth

OpenStudy (anonymous):

Idk

OpenStudy (dtan5457):

-164 divided by 2 is??

OpenStudy (jhannybean):

Just divide the first term by the second.

OpenStudy (anonymous):

\[ \frac{-164}{-82}=\ldots \]

OpenStudy (jhannybean):

Then the second term by the third, etc.

OpenStudy (mathmath333):

\(\Large a,ar,ar^2,ar^3...........\\ \Large -164, -82, -41, -20.5,..\\~\\~\\ \Large a=-164\\ \Large ar=-82\\~\\~\\ \Large \dfrac{ar}{a}=\dfrac{-82}{-164}\\ \Large r=\dfrac{82}{164}\\ \)

OpenStudy (anonymous):

lol wtf hacks?

OpenStudy (anonymous):

look at all the anonymous users LOLOLOL